# Differential Privacy for Deep Learning in Medicine

### v1

@@ Abstract
Differential privacy (DP) is a key technique for protecting sensitive patient data in medical deep learning (DL). As clinical models grow more data-dependent, balancing privacy with utility and fairness has become a critical challenge. This scoping review synthesizes recent developments in applying DP to medical DL, with a particular focus on DP-SGD and alternative mechanisms across centralized and federated settings. Using a structured search strategy, we identified 74 studies published up to March 2025. Our analysis spans diverse data modalities, training setups, and downstream tasks, and highlights the tradeoffs between privacy guarantees, model accuracy, and subgroup fairness. We find that while DP—especially at strong privacy budgets—can preserve performance in well-structured imaging tasks, severe degradation often occurs under strict privacy, particularly in underrepresented or complex modalities. Furthermore, privacy-induced performance gaps disproportionately affect demographic subgroups, with fairness impacts varying by data type and task. A small subset of studies explicitly addresses these tradeoffs through subgroup analysis or fairness metrics, but most omit them entirely. Beyond DP-SGD, emerging approaches leverage alternative mechanisms, generative models, and hybrid federated designs, though reporting remains inconsistent. We conclude by outlining key gaps in fairness auditing, standardization, and evaluation protocols, offering guidance for future work toward equitable and clinically robust privacy-preserving DL systems in medicine.
1 Introduction
@@
Deep learning (DL) has become a cornerstone of modern medical artificial intelligence (AI), driving advances in diagnostic imaging, survival prediction, biosignal analysis, and electronic health record (EHR) modeling [1, 2, 3]. However, the deployment of DL models in clinical settings raises critical concerns regarding patient privacy [4]. Medical data are inherently sensitive, and DL models trained on such data are vulnerable to various privacy attacks—including membership inference, model inversion, and identity disclosure—potentially exposing the presence or characteristics of individual patients in the training set [5, 6, 7, 8, 9].
@@
Differential privacy (DP) [10] provides formal guarantees to protect individuals within a dataset by bounding the influence of any single record on the model’s output. In DL, this is commonly implemented via differentially private stochastic gradient descent (DP-SGD) [11], which injects noise into clipped gradients during training (Fig. 1). Despite its theoretical strength, practical deployment of DP in healthcare DL introduces key trade-offs: noise reduces model utility and may disproportionately affect underrepresented subgroups, potentially reinforcing existing healthcare inequities [12, 13, 5].
@@
To better understand the state of DP in clinical DL, we conduct a comprehensive scoping review of empirical studies that apply DP mechanisms in medical contexts. While several prior reviews address DP broadly or focus on applications in generic domains, no existing synthesis systematically evaluates how DP interacts with clinical data, model design, privacy-utility trade-offs, fairness impacts, or adversarial robustness.
@@
This scoping review aims to critically synthesize current knowledge, identify key challenges and opportunities, and provide practical guidance for future research and clinical practice involving privacy-preserving DL models in medicine. In the following sections, we first review the applications of DP across clinical DL tasks, covering data modalities, training settings, and privacy accounting methods. We then analyze how architectural choices, normalization strategies, and pretraining affect the privacy-utility tradeoff, with a focus on model performance across varying privacy budgets. Next, we examine fairness impacts, highlighting subgroup disparities, the metrics used to evaluate them, and gaps in current practices. We also survey alternative privacy mechanisms beyond DP-SGD, including federated learning (FL), local perturbation, and hybrid methods. Finally, we synthesize empirical findings on privacy attacks and defenses, before concluding with a discussion of key limitations, open challenges, and recommendations for building robust, equitable, and privacy-preserving AI systems in medicine.
Figure 1: Overview of privacy risks and defenses in different phases of medical deep learning. (a) During model training, raw patient data—collected from clinical sources—can be vulnerable to leakage if accessed by unauthorized parties. Differential privacy is applied at this stage, typically by injecting noise into model gradients, to prevent individual data from being memorized. (b) After deployment, the trained model is exposed to potential inference attacks, such as membership inference, where adversaries aim to determine if specific individuals were in the training set. Models trained with differential privacy are more resistant to such attacks, enhancing patient confidentiality throughout the model lifecycle.
2 Methods
This scoping review was conducted in accordance with the Preferred Reporting Items for Systematic Reviews and Meta-Analyses extension for Scoping Reviews (PRISMA-ScR) guidelines [14, 15] to ensure methodological rigor and transparency [16]. The complete PRISMA-ScR checklist is provided in Supplementary Table 1. The methodological approach to data collection, article selection, and results synthesis is described in detail below.
2.1 Eligibility Criteria
2.2 Information Sources
We systematically searched four electronic databases: PubMed, IEEE Xplore, ACM Digital Library, and Web of Science. Database queries were performed on March 1, 2025, and search results were imported into Rayyan software [17], a tool designed to facilitate the systematic screening of literature and duplicate removal. A detailed description of the database-specific search queries is provided in Supplementary Table 2.
2.3 Search Strategy
The search strategy was carefully designed to identify publications at the intersection of DP, DL, and medical applications. Specifically, the search strings combined three primary keyword groups: (1) differential privacy techniques (e.g., “differential privacy,” “differentially private,” “DP-SGD”), (2) deep learning and related machine learning methodologies (e.g., “deep learning,” “machine learning,” “neural network,” “artificial intelligence”), and (3) healthcare and clinical applications (e.g., “medical,” “clinical,” “healthcare,” “biomedical,” “EHR,” “imaging,” “radiology,” “text,” “speech”).
2.4 Inclusion and Exclusion Criteria
The literature selection involved multiple screening steps: (1) Initial screening included automated duplicate removal and manual title and abstract screening in Rayyan. Articles were excluded if they lacked relevance to DP or healthcare applications. (2) Full-text reviews were conducted on the remaining articles to ensure alignment with inclusion criteria, specifically the empirical application of DP with DL techniques in medical contexts. Articles with purely theoretical content, lacking empirical evaluation, or applied to non-healthcare domains were excluded. The selection process ensured proportional representation across different medical application domains, including medical imaging, electronic health records, predictive analytics, and personalized medicine.
2.5 Synthesis of Results
The included studies were categorized using a structured two-step process. First, we grouped papers by their primary privacy mechanism, with a focus on applications of DP-SGD across clinical tasks and modalities, as summarized in Table 1. Design-related aspects—such as architectural choices, normalization strategies, and pretraining—are outlined in Table 2, while performance trends under varying privacy budgets are compiled in Table 3. Fairness evaluations, including subgroup analyses and reported metrics, are captured in Table 4. Studies employing alternative privacy mechanisms beyond DP-SGD are presented in Table 5, and empirical evaluations of attack resilience and defense strategies are detailed in Table 6.
3 Included studies
@@
A total of 4,242 records were retrieved through systematic searches across four databases: PubMed, IEEE Xplore, ACM Digital Library, and Web of Science. After automated and manual removal of duplicates, 2,600 unique records remained. During the initial title and abstract screening phase, 2,529 articles were excluded based on the predefined inclusion and exclusion criteria, primarily due to lack of relevance to medical applications or absence of DP components.
The remaining 71 articles underwent full-text screening to assess their eligibility and methodological alignment with the review objectives. At this stage, additional exclusions were made based on insufficient empirical content, non-healthcare focus, or redundancy in covered topics, resulting in the exclusion of 20 more papers.
Ultimately, 51 studies met all criteria and were included from the database search. To supplement this, 23 additional relevant studies were identified through manual searching of reference lists and preprint repositories, ensuring coverage of the most recent and impactful contributions. In total, 74 papers were included in this scoping review. An overview of the entire screening and selection process is illustrated in Fig. 2. Additional descriptive characteristics of DP publications in medical AI are shown in Fig. 3.
@@
Figure 2: Summary of study inclusion and descriptive statistics. (a) PRISMA flow diagram illustrating the study selection process for this scoping review. The diagram details the number of records identified through database and manual searches, duplicates removed, records screened by title and abstract, full-text articles assessed for eligibility, and the final number of studies included. (b) Publications per year on DP applications in medical DL, as of March 1, 2025. (c) Geographic distribution of included studies based on author affiliations; circle size reflects the number of publications with at least one affiliated author per country. Country outlines are based on Natural Earth.
Figure 3: Overview of study characteristics across included papers. Distribution of papers grouped by (a) modality or data type, (b) downstream prediction task, and (c) utilized neural network architectures. CNN: convolutional neural network; ECG: electrocardiogram; EEG: electroencephalography; EHR: electronic health record; GAN: generative adversarial network; GNN: graph neural network; MLP: multilayer perceptron; ViT: vision transformer.
4 Theoretical Background
We begin by establishing the core concepts and notation that form the basis of our discussion on DP in DL, followed by a detailed exposition of relevant privacy mechanisms and their integration into DL workflows.
Throughout this paper, we use uppercase bold letters (e.g., $\mathbf{X}$) for matrices, lowercase bold letters (e.g., $\mathbf{w}$) for vectors, uppercase letters (e.g., $D$) for datasets, and lowercase letters (e.g., $x$) for scalar variables or individual data samples.
4.1 Differential Privacy Definitions
@@
DP was introduced by Dwork et al. [18] as a mathematically rigorous framework to limit the influence of any single individual’s data on the output of a computation. This protection is achieved by injecting controlled noise into a computation’s output, thereby masking the presence or absence of individual records.
@@
Here, $\varepsilon$ controls the worst-case multiplicative difference in output probabilities between $D$ and $D^{\prime}$, while $\delta$ allows for a small additive probability of failure. When $\delta=0$, we recover the original, stricter definition of $\varepsilon$-DP.
In practice, $\delta$ is often set inversely proportional to dataset size ($\delta\sim 1/n$), ensuring that the likelihood of privacy failure remains negligible. Together, the $(\varepsilon,\delta)$ pair quantifies the privacy guarantee, with smaller values implying stronger protection.
Rényi differential privacy (RDP) [19] is a generalization of DP that leverages Rényi divergence [20] to provide tighter composition guarantees. The Rényi divergence of order $\alpha>1$ between the output distributions of $M$ on $D$ and $D^{\prime}$ is defined as:
@@
The amount of noise required to achieve DP depends on the sensitivity of the function $f$ being computed, which quantifies how much $f$ can change when a single input is modified.
For vector-valued outputs, the $\ell_{2}$-sensitivity of $f$ is:
@@
DP mechanisms can be applied at various stages of the learning pipeline, including the input [21, 22, 23], training process [11, 24], and model output [25, 26]. Among these, training-time DP has been shown to yield better accuracy and robustness [27].
4.2 Deep Learning and Optimization via SGD
DL refers to models that learn hierarchical representations from data via neural networks with multiple layers [28, 29]. Training these models involves minimizing an empirical risk objective:
@@ 4.2 Deep Learning and Optimization via SGD
where $L$ is the loss function and $\boldsymbol{\theta}$ the model parameters. This is typically optimized using stochastic gradient descent (SGD).
Variants of SGD differ by the batch size used in each update. In practice, mini-batch SGD is most common:
@@
4.3 Applying Privacy in Deep Neural Networks
Deep neural networks are vulnerable to privacy attacks targeting the input data, gradients, or model parameters [30]. Gradients are particularly sensitive, especially in federated settings where they are transmitted across devices [31]. Since gradients are computed directly from the input data, they can leak private information under adversarial scrutiny.
A common defense is to inject noise into gradients during training. DP-SGD [11] achieves this by modifying the gradient computation at each iteration. For each example $i$ in a mini-batch $\mathcal{B}$:
@@ 4.3 Applying Privacy in Deep Neural Networks
Here, $C$ is the clipping norm and $\sigma$ is the noise multiplier. Fig. 4 illustrates this process. Clipping limits any individual’s contribution, while the added noise ensures DP guarantees.
Variants such as dynamic DP-SGD [32] adapt $C$ and $\sigma$ over time to improve utility during training. Other refinements target sparse gradients, which arise in applications like language models. A recent method [33] improves DP-SGD efficiency by focusing on non-zero gradient components, offering better performance in high-dimensional or sparse settings.
Figure 4: Overview of the DP-SGD training process. (a) In each iteration, a mini-batch of data is sampled. (b) Per-sample gradients are computed and clipped to a fixed norm to bound sensitivity. (c) Noise—typically Gaussian—is added to the aggregated clipped gradients to enforce differential privacy. (d) The model is updated using the noisy gradients. This process is repeated over multiple iterations, yielding a trained model with formal privacy guarantees.
4.4 Private Aggregation of Teacher Ensembles
Private aggregation of teacher ensembles (PATE), introduced by Papernot et al. [34], is a differentially private learning framework designed to protect sensitive training data by aggregating predictions from multiple independently trained models, referred to as ”teachers.” Each teacher model $T_{k}$ is trained on a disjoint subset $D_{k}$ of a sensitive dataset:
$D=\bigcup_{k=1}^{K}D_{k},\quad D_{i}\cap D_{j}=\emptyset\text{ for }i\neq j$ (17)
Given an input $x$, each teacher $T_{k}$ produces a class prediction $T_{k}(x)\in\{1,\dots,C\}$, and votes are tallied across the ensemble:
$v_{c}(x)=\sum_{k=1}^{K}\mathbb{I}[T_{k}(x)=c]$ (18)
To preserve privacy, Laplace noise is added to the vote counts:
$\tilde{v}_{c}(x)=v_{c}(x)+\text{Lap}(1/\lambda)$ (19)
Here, $\lambda$ controls the privacy level, with smaller values inducing stronger noise and stronger privacy.
A student model $S$ is then trained using public unlabeled data $D_{\text{public}}$, where noisy labels are obtained by querying the teacher ensemble. The final label assigned to each sample is:
$\hat{y}(x)=\arg\max_{c}\tilde{v}_{c}(x)$ (20)
The student is trained to minimize a supervised loss using the noisy labels:
$\mathcal{L}_{\text{student}}=\frac{1}{N}\sum_{i=1}^{N}\ell(S(x_{i}),\hat{y}(x_{i}))$ (21)
The total privacy loss is tracked using a privacy accountant, such as the moments accountant [11], ensuring that the cumulative privacy budget $\epsilon$ is not exceeded.
Compared to DP-SGD, which perturbs gradients within a single model during training, PATE decouples private learning from student training by using noise-perturbed knowledge transfer. PATE generally performs well in semi-supervised settings or when class imbalance can be mitigated via teacher ensembling. For example, a comparative study [35] showed that PATE-based GAN training often yields better utility-privacy tradeoffs than DP-SGD in cases where sensitive training data can be partitioned effectively.
4.5 Alternative Methods for Privacy Preservation
Beyond DP-SGD and PATE, several additional strategies have been developed to support privacy-preserving DL. One widely adopted framework is federated learning (FL), which enables collaborative model training across decentralized institutions or devices without requiring raw data to leave local storage [31]. FL can be implemented in centralized configurations—where local updates are sent to a coordinating server—or in decentralized topologies, where clients communicate directly with peers based on a predefined graph structure [36]. FL is increasingly used in clinical AI [4, 37, 38, 39], and is often combined with DP by adding noise to model updates. In such settings, privacy can be ensured either through centralized mechanisms—where noise is added after server aggregation—or local mechanisms, where clients inject noise prior to communication [40, 41].
Another major line of research involves secure multiparty computation, originally introduced by Yao [42]. It enables multiple parties to jointly compute functions over private inputs without revealing those inputs to each other. In DL, secure multiparty computation has been used to train or evaluate models securely across institutional boundaries [43]. Some recent works combine secure multiparty computation with DP [44, 45], adding noise either at each training iteration or once at the end of training. These hybrid designs differ in trade-offs: for instance, [44] focuses on communication efficiency through incremental protection, while [45] emphasizes accuracy preservation by applying noise post hoc.
Several methods aim to protect privacy by modifying gradients during training. These include clipping, pruning, masking, or injecting noise into gradients to obscure sensitive signals [46]. While some evidence suggests that such gradient masking can limit leakage without harming utility, other studies have shown that these defenses can be circumvented by more sophisticated adversaries [47].
An alternative line of defense involves perturbing internal feature representations rather than gradients. These feature-level defenses add noise or transformations at intermediate network layers, reducing the risk of inversion or memorization [48, 49, 50].
Synthetic data generation offers another privacy-preserving strategy. In this approach, a generative model (e.g., GAN or diffusion model) is trained on private data—sometimes under DP constraints—and is then used to produce synthetic datasets for downstream model training [51, 52]. This decouples the final model from direct exposure to sensitive records, reducing re-identification risk.
Classical data anonymization methods are still used in some clinical contexts. These include $k$-anonymity [53], $l$-diversity [54], and $t$-closeness [55], which aim to ensure that individual records are indistinguishable within equivalence classes. However, these approaches lack formal guarantees and are often vulnerable to auxiliary information attacks.
Another direction of research considers how architectural choices impact privacy. For example, empirical studies suggest that transformer-based models tend to leak more information than convolutional neural networks (CNNs), possibly due to higher capacity and attention mechanisms [56]. Components such as normalization layers, activation functions, and attention structures can influence a model’s vulnerability to privacy attacks.
Finally, cryptographic tools such as homomorphic encryption and secure aggregation are increasingly applied to protect data during training and communication [57]. In FL, for instance, gradients can be encrypted locally before transmission, adding an additional layer of protection against malicious servers or eavesdroppers.
These alternative mechanisms illustrate the diversity of tools available for protecting patient data in clinical machine learning. While our review focuses primarily on DP-SGD and PATE due to their theoretical foundations and empirical uptake, the broader ecosystem of privacy-preserving techniques—including encryption, secure computation, gradient manipulation, and architectural design—offers complementary strategies that may be better suited for specific tasks or deployment environments.
5 Applications of Differential Privacy in Clinical Deep Learning
5.1 Overview of Application Domains
@@
DP has been applied to a wide range of clinical use cases, spanning diverse data modalities, deployment settings, and downstream tasks. Table 1 provides a summary of studies that applied DP via DP-SGD, the most widely adopted method in this domain. These studies encompass both centralized and FL settings, and cover modalities such as medical imaging (e.g., X-ray, CT, MRI), tabular data (e.g., EHR), biosignals (e.g., EEG, ECG), and multimodal inputs.
@@
Adnan et al. [58] Histopathology FL Classification (lung cancer) GM/RDP
Al Aziz et al. [59] EHR Centralized EHR generation GM/MA
Chilukoti et al. [60] Histopathology Centralized Gastric cancer classification GM/RDP
Chin-Cheong et al. [61] EHR Centralized EHR generation GM/MA
Daum et al. [62] Cardiac cine MRI (4D) Centralized 3D MRI generation GM/RDP
DP-LPR [63] Genomic (DNA-SNPs) Centralized Gene interaction detection LM/NA
DP-SSLoRA [64] Chest X-ray Centralized Classification GM/RDP
DPFedSAM-Meas [65] Chest X-ray FL Classification GM, LM/RDP
Fan et al.[66] Tabular data Centralized Regression (survival analysis) GM/RDP
FedGAN [67] Chest X-ray FL Classification GM/NA
Guo et al. [68] Dermoscopic photography Centralized Classification GM/NA
Hatamizadeh et al. [69] Chest X-ray, fundoscopy FL Classification GM/NA

### v2

# Differential privacy for medical deep learning: methods, tradeoffs, and deployment implications

@@ Abstract
Differential privacy (DP) is a key technique for protecting sensitive patient data in medical deep learning (DL). As clinical models grow more data-dependent, balancing privacy with utility and fairness has become a critical challenge. This scoping review synthesizes recent developments in applying DP to medical DL, with a particular focus on DP-SGD and alternative mechanisms across centralized and federated settings. Using a structured search strategy, we identified 74 studies published up to March 2025. Our analysis spans diverse data modalities, training setups, and downstream tasks, and highlights the tradeoffs between privacy guarantees, model accuracy, and subgroup fairness. We find that while DP, especially at strong privacy budgets ($\epsilon\approx$10), can preserve performance in well-structured imaging tasks, severe degradation often occurs under strict privacy ($\epsilon\approx$1), particularly in underrepresented or complex modalities. Furthermore, privacy-induced performance gaps disproportionately affect demographic subgroups, with fairness impacts varying by data type and task. A small subset of studies explicitly addresses these tradeoffs through subgroup analysis or fairness metrics, but most omit them entirely. Beyond DP-SGD, emerging approaches leverage alternative mechanisms, generative models, and hybrid federated designs, though reporting remains inconsistent. We conclude by outlining key gaps in fairness auditing, standardization, and evaluation protocols, offering guidance for future work toward equitable and clinically robust privacy-preserving DL systems in medicine.
Introduction
@@
Deep learning (DL) has become a cornerstone of modern medical artificial intelligence (AI), driving advances in diagnostic imaging, survival prediction, biosignal analysis, and electronic health record (EHR) modeling [rajpurkar2022ai, tayebi2024large, haug2023artificial, tayebi2024treasure]. However, the deployment of DL models in clinical settings raises critical concerns regarding patient privacy [66kaissis2021end]. Medical data are inherently sensitive, and DL models trained on such data are vulnerable to various privacy attacks, including membership inference, model inversion, and identity disclosure, potentially exposing the presence or characteristics of individual patients in the training set [22tayebi2024preserving, usynin2021adversarial, wang2021variational, haim2022reconstructing, carlini2023extracting].
@@
Differential privacy (DP) [dwork2014algorithmic] provides formal guarantees to protect individuals within a dataset by bounding the influence of any single record on the model’s output. In DL, this is commonly implemented via differentially private stochastic gradient descent (DP-SGD) [DPSGD], which injects noise into clipped gradients during training (Fig. 1). Despite its theoretical strength, practical deployment of DP in healthcare DL introduces key tradeoffs: noise reduces model utility and may disproportionately affect underrepresented subgroups, potentially reinforcing existing healthcare inequities [tran2021decision, 13TayebiDomainTransfer, 22tayebi2024preserving].
@@
To better understand the state of DP in clinical DL, we conduct a comprehensive scoping review of empirical studies that apply DP mechanisms in medical contexts. While several prior reviews, such as in [zhao2019differential, demelius2025recent, fu2024differentially, liu2024survey], address DP broadly or focus on applications in generic domains, no existing synthesis systematically evaluates how DP interacts with clinical data, model design, privacy-utility tradeoffs, fairness impacts, or adversarial robustness.
@@
This scoping review aims to critically synthesize current knowledge, identify key challenges and opportunities, and provide practical guidance for future research and clinical practice involving privacy-preserving DL models in medicine. Throughout this review, ”training” refers to model optimization on available data; ”validation” refers to performance evaluation on a held-out subset of the same dataset; and ”external validation” denotes evaluation on data from a different site, institution, or patient population. ”Deployment” refers to releasing a model for operational clinical use (e.g., inside a hospital system or federated network). ”Monitoring” refers to post-deployment performance tracking, including detection of performance degradation or subgroup fairness drift over time. In federated learning (FL) [FederatedLearningSource], ”clients” denote data-holding sites; ”centralized” refers to training on pooled data. We use “DP-SGD,” “local DP,” “privacy budget,” and “privacy accountant” with their standard meanings in DP literature. In the following sections, we first review the applications of DP across clinical DL tasks, covering data modalities, training settings, and privacy accounting methods. We then analyze how architectural choices, normalization strategies, and pretraining affect the privacy-utility tradeoff, with a focus on model performance across varying privacy budgets. Next, we examine fairness impacts, highlighting subgroup disparities, the metrics used to evaluate them, and gaps in current practices. We also survey alternative privacy mechanisms beyond DP-SGD, including FL, local perturbation, and hybrid methods. Finally, we synthesize empirical findings on privacy attacks and defenses, before concluding with a discussion of key limitations, open challenges, and recommendations for building robust, equitable, and privacy-preserving AI systems in medicine.
Figure 1: Overview of differential privacy (DP) applied to model training and inference. (a) Training samples are processed individually to compute per-sample gradients, which are clipped to limit sensitivity and then perturbed by noise sampled from a Gaussian distribution with variance $\sigma^{2}$. Higher noise (larger $\sigma$) yields stronger privacy (smaller $\epsilon$), and the resulting noisy gradients are aggregated to update model parameters. (b) DP bounds how much the model’s output distribution can change when a single individual’s data is added or removed from the training set. When noise is low (larger $\epsilon$), outputs for neighboring datasets may differ noticeably, making membership inference easier. With stronger privacy (smaller $\epsilon$), output distributions overlap and individual influence becomes indistinguishable. The parameter $\delta$ represents the probability that the DP guarantee may not hold. (c) In the absence of DP, an attacker with access to a deployed model may exploit output differences to infer whether a specific sample was included in training, potentially revealing sensitive patient information (e.g., metadata or identifiers). Representative chest X-ray images are provided by the ChestX-ray14 dataset from NIH Clinical Center [wang2017chestx].
Results
@@
A total of 4,242 records were retrieved through systematic searches across PubMed, IEEE Xplore, ACM Digital Library, and Web of Science. After automated and manual duplicate removal, 2,600 unique records remained. During title and abstract screening, 2,529 records were excluded based on predefined eligibility criteria, primarily due to lack of medical relevance or absence of differential privacy components. The remaining 71 articles underwent full-text assessment, where 20 additional papers were excluded due to insufficient empirical content or misalignment with the review scope. Ultimately, 51 studies were retained from the database search, and an additional 23 studies were identified through citation tracking and preprints, resulting in a total of 74 included papers. The full screening and selection workflow is shown in Fig. 2, with descriptive publication trends summarized in Fig. 3.
@@
Figure 2: Preferred reporting items for systematic reviews and meta-analyses extension for scoping reviews (PRISMA-ScR) flow diagram illustrating the selection process for this study. The diagram [tricco2018prisma, page2021prisma] details the number of records identified through database and manual searches, duplicates removed, records screened by title and abstract, full-text articles assessed for eligibility, and the final number of studies included.
To structure the analysis, the included studies were first organized according to their primary privacy mechanism. Applications of DP-SGD, the most widely used approach in the field [66kaissis2021end, 22tayebi2024preserving, 67ziller2024reconciling, de2022unlocking, chua2024private], are summarized in Table 1. Design factors influencing the privacy–utility tradeoff, including model architectures, normalization strategies, and pretraining choices, are presented in Table 2, while performance trends across different privacy budgets are compiled in Table 3. Fairness evaluations and subgroup analyses are captured in Table 4. Studies using alternative privacy mechanisms beyond DP-SGD are listed in Table 5, and evaluations of attack resilience are shown in Table 6.
Figure 3: Overview of study characteristics across included papers. (a) Geographic distribution of included studies based on author affiliations. Color intensity reflects the number of publications with at least one affiliated author per country; all affiliations of all the coauthors of each study are considered in this statistic. The top 4 contributing countries include the USA, China, the UK, and Germany, with 29, 24, 13, and 12 studies, respectively. Distributions of included studies (n=74) are shown with bar diagrams by (b) publication year, (c) publication type, and (d) training paradigm. Distributions of DP-SGD studies (n=67) are shown with bar diagrams by (e) data modality, (f) downstream tasks, and (g) architecture type. Some studies used multiple modalities, downstream tasks, or architectures; all such instances were counted, so categories are not mutually exclusive. Bar heights indicate counts. FL: federated learning; MLP: multilayer perceptron; CNN: convolution neural networks; GAN: generative adversarial network; GNN: graph neural network; MLP: multilayer perceptron; ViT: vision transformer.
Theoretical background
Before presenting the empirical findings, we provide the necessary theoretical background to clarify how DP is mathematically defined and how these guarantees extend to DL workflows. This section explains the formal privacy parameters $(\epsilon,\delta)$, their composition when data or models are accessed multiple times, and how DP is operationalized in DL through gradient clipping and noise addition.
@@
DP was introduced by Dwork et al. [TheAlgorithmicFoundationsofDifferentialPrivacy] as a mathematically rigorous framework to limit the influence of any single individual’s data on the output of a computation. This protection is achieved by injecting controlled noise into a computation’s output, thereby masking the presence or absence of individual records.
@@
Here, $\varepsilon$ controls the worst-case multiplicative difference in output probabilities between $D$ and $D^{\prime}$, while $\delta$ allows for a small additive probability of failure. When $\delta=0$, we recover the original, stricter definition of $\varepsilon$-DP. In practice, $\delta$ is often set inversely proportional to dataset size ($\delta\sim 1/n$), ensuring that the likelihood of privacy failure remains negligible. Together, the $(\varepsilon,\delta)$ pair quantifies the privacy guarantee, with smaller values implying stronger protection.
Rényi differential privacy (RDP) [RDP] is a generalization of DP that leverages Rényi divergence [RD] to provide tighter composition guarantees. The Rényi divergence of order $\alpha>1$ between the output distributions of $M$ on $D$ and $D^{\prime}$ is defined as:
@@
In real-world deployments, differentially private mechanisms are rarely applied only once [67ziller2024reconciling, kaissis2020secure]. Instead, the same dataset may be accessed repeatedly, for example, during iterative model training (such as DP-SGD, which performs many gradient updates), or after deployment through user queries. DP therefore requires tracking the cumulative privacy loss across all uses of the data [DPSGD]. The sequential composition property states that if two differentially private mechanisms are applied to the same dataset, with privacy parameters $(\varepsilon_{1},\delta_{1})$ and $(\varepsilon_{2},\delta_{2})$, then releasing the outputs of both mechanisms provides $(\varepsilon_{1}+\varepsilon_{2},\delta_{1}+\delta_{2})$-DP. Thus, each access to the data consumes part of the overall privacy budget, and the guarantee becomes weaker as more outputs are released [dwork2014algorithmic].
For iterative methods such as DP-SGD, naïve linear composition can substantially overestimate the privacy cost. RDP enables tighter accounting of cumulative privacy loss across many iterations, resulting in improved bounds such as:
$\varepsilon_{\text{total}}\approx\sqrt{T}\cdot\varepsilon_{\text{per iteration}}$ (4)
where $T$ is the number of training iterations. Modern DP libraries, including Opacus [opacus_neurips2021], implement privacy accountants based on this formulation and stop training once a predefined privacy budget is exhausted.
Repeated use also introduces additional considerations once a private model is deployed. Multiple users may query the same model, and each query consumes privacy budget from the same underlying dataset. The DP guarantee assumes that users act independently; if users share outputs or coordinate queries, their combined set of queries must be treated as a single adversary, and privacy loss accumulates accordingly. In practice, complete non-collusion cannot be guaranteed, especially in clinical environments where multiple analysts or systems may interact with the same model. Therefore, deployed DP systems incorporate operational safeguards, such as authentication, per-user privacy budget allocation, query logging, audit trails, and automatic disabling of access once the available privacy budget is exhausted [cormode2018privacy]. Privacy budget tracking is increasingly required in regulated settings (e.g., hospital data governance platforms), reflecting that DP is not only a training-time guarantee but a limited resource that must be monitored throughout deployment [66kaissis2021end].
The amount of noise required to achieve DP depends on the sensitivity of the function $f$ being computed, which quantifies how much $f$ can change when a single input is modified. For vector-valued outputs, the $\ell_{2}$-sensitivity of $f$ is:
@@
DP mechanisms can be applied at various stages of the learning pipeline, including the input [noise_at_input_1, noise_at_input_2, noise_at_input_3], training process [DPSGD, noise_during_training_label], and model output [noise_at_output_1, noise_at_output_2]. Among these, training-time DP has been shown to yield better accuracy and robustness [trade-offs].
DL refers to models that learn hierarchical representations from data via neural networks with multiple layers [DeepLearningLeCun, DeepLearningSchmidhuber]. Training these models involves minimizing an empirical risk objective:
@@
where $L$ is the loss function and $\boldsymbol{\theta}$ the model parameters. This is typically optimized using stochastic gradient descent (SGD). Variants of SGD differ by the batch size used in each update. In practice, mini-batch SGD is most common:
@@
Deep neural networks are vulnerable to privacy attacks targeting the input data, gradients, or model parameters [ASurveyofPrivacyAttacksinMachineLearning]. Gradients are particularly sensitive, especially in federated settings where they are transmitted across devices [FederatedLearningSource]. Since gradients are computed directly from the input data, they can leak private information under adversarial scrutiny. A common defense is to inject noise into gradients during training. DP-SGD [DPSGD] achieves this by modifying the gradient computation at each iteration. For each example $i$ in a mini-batch $\mathcal{B}$:
@@
Here, $C$ is the clipping norm and $\sigma$ is the noise multiplier. Clipping limits any individual’s contribution, while the added noise ensures DP guarantees.
Overview of applications of DP in clinical DL
@@
DP has been applied to a wide range of clinical use cases, spanning diverse data modalities, deployment settings, and downstream tasks. Table 1 provides a summary of studies that applied DP via DP-SGD, the most widely adopted method in this domain [DPSGD, de2022unlocking, chua2024private]. These studies encompass both centralized and FL settings, and cover modalities such as medical imaging (e.g., X-ray, CT, MRI), tabular data (e.g., EHR), biosignals (e.g., EEG, ECG), and multimodal inputs (Fig. 4).
@@
Adnan et al. [36adnan2022federated] Histopathology images FL Classification GM/RDP
Al Aziz et al. [84DifferentiallyPrivateMedicalTextsGenerationUsingGenerativeNeuralNetworks] EHR Centralized Tabular data generation GM/MA
Chilukoti et al. [86chilukoti2025differentiallyprivatefinetunednfnet] Histopathology images Centralized Classification GM/RDP
Chin-Cheong et al. [74chin2020generation] EHR Centralized Tabular data generation GM/MA
Daum et al. [71OnDifferentiallyPrivate3DMedicalImageSynthesiswithControllableLatentDiffusionModels] Cardiac cine MRI Centralized Image generation GM/RDP
DP-LPR [25zhang2022secure] Genomic data Centralized Feature selection LM/NA
DP-SSLoRA [24DP-SSLoRA] Chest X-ray Centralized Classification GM/RDP
DPFedSAM-Meas [50DPFedSAM-Meas] Chest X-ray FL Classification GM, LM/RDP
Fan et al.[26fan2023mitigating] Tabular data Centralized Regression GM/RDP
FedGAN [30nguyen2021federated] Chest X-ray FL Classification GM/NA
Guo et al. [53Adifferentialprivacybasedprototypicalnetworkformedicaldatalearning] Dermoscopic photography Centralized Classification GM/NA
Hatamizadeh et al. [33hatamizadeh2023gradient] Chest X-ray, fundoscopy FL Classification GM/NA